OpenStudy (abbycross167):

Can someone please check my work? Question~The product of a number and -5 less than 7.2 My Answer~ X*7.2-(-5) What my teacher said~ 7.2 - (-5x) simplifying is? ~What does my teacher mean by that? Do I need to add "7.2 - (-5x)" together to get 12.2 as the answer?

jimthompson5910 (jim_thompson5910):

`The product of a number and -5` means we have some unknown number x times -5 which is equal to -5*x or just -5x

jimthompson5910 (jim_thompson5910):

since we're saying "less than 7.2" this means that we're subtracting whatever the result of -5x is from 7.2, so we say 7.2 - (-5x)

jimthompson5910 (jim_thompson5910):

whenever you subtract a negative, you're really adding a positive 7.2 - (-5x) = 7.2 + 5x we can't go any further because 7.2 and 5x aren't like terms. We cannot combine them

jimthompson5910 (jim_thompson5910):

we can't say 7.2 + 5x = 12.2x or 7.2+5x = 12.2 (both equations are false)

OpenStudy (abbycross167):

Oh okay, well what did she mean by "simplifying is"?

jimthompson5910 (jim_thompson5910):

just turning `7.2 - (-5x)` into `7.2 + 5x`

OpenStudy (abbycross167):

oh okay, Thank you very much @jim_thompson5910

OpenStudy (ghostgate):

I'll explain it this way: The product of a number and -5 less than 7.2 "The product" = Multiplication "A number" = x "and -5" = Concludes that the number is multiplied against -5, giving you -5x. "less than" = Subtraction however words it in a tricky manner, making you finalize with: -5x - 7.2 - Even though it would be this: 7.2 - (-5x) or 7.2 + 5x

OpenStudy (abbycross167):

Thank you too! @Ghostgate

jimthompson5910 (jim_thompson5910):

np

OpenStudy (ghostgate):

Yeah, I was a bit late, but you're welcome! ;)